The activation energy Ea for a particular reaction is 50.0 kJ/mol. How much faster is the reaction at 313 K than at 310 K? (R = 8.314 J/mol • K) A) 0.185 B) 1.20 C)2.00 x 10^5 D) 1.002 E) 0.83
Answer #1

log (K2/K1) = Ea / 2.303 R [1/T1 - 1/T2]

T1 = 310 K , T2 = 313 K

log (K2/K1) = (50 x 1000 / 8.314 x 2.303) [1/310 - 1/313]

log (K2/K1) = (50000 / 19.147) [0.00322 - 0.00319]

log (K2/K1) = 50000 / 19.147 [0.00003]

log (K2/K1) = 0.0783

K2/K1 = 1.20

answer = option B = 1.20